1) Describe the expected changes in bond strength when NO is reduced to NO- . Explain (5 pts).

2) Is this process predicted to be favorable or unfavorable? Explain. (5 pts)
Answer #1

1) Electronic configuration of NO is

Bond order = )Electrons in bonding orbital - Electrons on antibonding orbital)/2 = (10 -5)/2 = 2.5

Now, electronic configuration of NO- is

Bond order = (10 -6)/2 = 2.

Now, Bond order Bond strength.

As bond order of NO > NO- .

Therefore bond strength of NO is higher than NO-.

2) Antibonding orbitals has energy than the energy of atomic orbitals. Electrons in this orbital lead to repulsion between atoms. Antibonding orbitals are less stable than bonding orbitals.

In the transformation of NO to NO- the incoming electron is added to the antibonding orbital then it requires energy. Hence the reduction of NO to NO- is unfavourable.